# GraSP: A Unified Graph-Based Framework for Scalable Generation, Quality Tagging, and Management of Synthetic Data for SFT and DPO

### v1

@@ Abstract
The advancement of large language models (LLMs) is critically dependent on the availability of high-quality datasets for Supervised Fine-Tuning (SFT), alignment tasks like Direct Preference Optimization (DPO), etc. In this work, we present a comprehensive synthetic data generation framework that facilitates scalable, configurable, and high-fidelity generation of synthetic data tailored for these training paradigms. Our approach employs a modular and configuration-based pipeline capable of modeling complex dialogue flows with minimal manual intervention. This framework uses a dual-stage quality tagging mechanism, combining heuristic rules and LLM-based evaluations, to automatically filter and score data extracted from OASST-formatted conversations, ensuring the curation of high-quality dialogue samples. The resulting datasets are structured under a flexible schema supporting both SFT and DPO use cases, enabling seamless integration into diverse training workflows. Together, these innovations offer a robust solution for generating and managing synthetic conversational data at scale, significantly reducing the overhead of data preparation in LLM training pipelines.
@@ 1 Introduction
To address this, we introduce GraSP, a general-purpose framework for scalable synthetic data generation. GraSP combines low-code, YAML-based configuration with modular, graph-driven orchestration to support complex workflows with branching, looping, and conditionals. It enables the reuse of graphs as subgraphs, ensures reliable execution through integrated validation and checkpointing, and natively supports multimodal inputs and agent based data generation. Additionally, GraSP offers unified dataset I/O across HuggingFace and local formats, supports quality tagging, and produces outputs compatible with OASST-style formatting for seamless downstream use.
@@ 2 Related Work
Table 1: Comparison of GraSP with popular frameworks across key capabilities.
Category Feature GraSP Distilabel SDG Curator Synthetic Data Kit
@@
In summary, while existing frameworks each offer valuable features for synthetic data generation, they typically address isolated aspects of the broader workflow. GraSP stands out by providing a unified, extensible approach that brings together the critical capabilities needed for modern, complex, and multimodal data generation pipelines.
3 GraSP Framework
GraSP is a modular and extensible system designed for large-scale, programmable data generation. It supports configurable orchestration through a graph abstraction that enables reusable, auditable, and resumable workflows. The framework is designed for both research and production pipelines, with pluggable model backends and modular task authoring support.
Figure 1: High-level GraSP architecture.
@@ 3.1 System Architecture
GraSP is guided by three principles—Scalability (streaming data sources, resumable jobs, JSONL/Parquet/HF outputs), Modularity (YAML-defined DAG workflows with conditional logic), and Reusability (versioned, reusable graphs, nodes, and validators). Figure 1 shows its core components:
@@ 3.2 Pipeline Components
GraSP pipelines are defined declaratively in YAML, promoting low-code, reproducible workflow construction. Each pipeline consists of three configuration blocks:
@@
Specifies input and output sources, format handling (CSV, JSONL, Parquet), streaming options, and inline preprocessing (e.g., renaming, filtering, combining). Supports both data-backed and data-less generation scenarios.
@@
Defines a DAG of computational nodes, including LLM calls, lambdas, agents, or subgraphs. Supports conditional branching, step-wise pre/post-processing, and is compiled to a LangGraph-compatible representation.
@@
Refer to Appendix A for detailed configuration options and schema definitions.
@@ 3.3 Key Features
GraSP brings together robust design abstractions and practical scalability for real-world use cases. Specifically, our contributions include:
@@
    Low-Code, Modular Graph Configuration: GraSP combines a YAML-based interface with LangGraph-style agents and a custom DAG engine, enabling concise, extensible definitions of complex workflows with branching, looping, and conditionals. B
@@
    Reusable Recipes (Subgraphs): This feature enables us to use common graph components which can be reused across tasks, promoting modularity. For instance, the Evolve INSTRUCT recipe (Figure  2) encapsulates a modular subgraph that receives seed instructions and applies either depth-based or breadth-based evolution strategies via a routing node (Strategy) [7]. This subgraph can be invoked repeatedly across different flows, enhancing composability and reducing redundancy.
    Figure 2: Instruction evolution subgraph and judgment loop used within GraSP pipelines.
@@
    Multimodal Support: GraSP extends beyond text-only workflows by natively handling audio and image inputs alongside text. Through unified I/O adapters, it transparently loads local or remote media in various formats, encodes them as base64 data URLs for LLM API compatibility, and supports multiple media fields per record. This enables workflows for tasks such as speech recognition, audio classification, document analysis, and visual QA. Round-tripping ensures outputs can be saved back into HuggingFace datasets in their original formats for reproducibility and downstream use.
@@
    Agentic Execution: GraSP enables the creation of autonomous, tool-using agents built on the ReAct reasoning-and-acting paradigm via LangGraph. Agent nodes extend LLM nodes with capabilities for dynamic tool invocation, multi-turn reasoning, and conditional decision-making. Developers can specify a library of callable tools, inject context-specific system messages at arbitrary conversation turns, and configure pre/post-processing hooks for fine-grained control over input and output. This allows pipelines to handle exploratory tasks, iterative search, and interactive decision flows in a modular, low-code manner.
@@
    Structured Output Generation: GraSP provides a flexible framework for generating and validating structured outputs from LLMs, reducing post-processing effort and ensuring reliable formats. It supports both class-based schemas (via Pydantic) and YAML-defined schemas, with automatic type handling and optional custom validation rules. Structured output generation works natively with OpenAI and vLLM models, and falls back to JSON schema validation for other backends. This allows developers to define precise field types, attach descriptions, and enforce constraints directly at generation time.
@@
    Resumability: GraSP supports fault-tolerant, restartable execution of long-running jobs. In the event of a failure, execution can gracefully shut down and later resume from the last recorded checkpoint without reprocessing completed steps. This is particularly valuable for large-scale or streaming workloads where partial progress should be preserved. Checkpoints store both intermediate outputs and node-level metadata, enabling accurate restoration of execution state.
@@ 4 Results and Impact
Using GraSP’s end-to-end graph-driven pipelines, we have generated billions of tokens of synthetic SFT, DPO, and CPT data, and hundreds of thousands of domain-specific records, a few of them are summarized in Table.
Throughput & Latency: Baseline vs. GraSP Orchestration
To illustrate GraSP’s impact on end-to-end processing time, we compared two representative code-generation tasks running over 1,000 records using a naive, single-threaded pipeline (baseline) versus GraSP’s parallel, non-blocking orchestration. As shown in Table 2, GraSP delivers a 3×–4× reduction in wall-clock time.
Task Baseline GraSP Speedup (×)
Coding Question Grading 2054.7 624.7 $\approx 3.3$
Glaive Code Assistant 2270.6 572.4 $\approx 4.0$
Table 2: End-to-End Processing Times (in seconds): Baseline vs. GraSP (1,000 records)
By replacing a sequential loop with GraSP’s asynchronous agent scheduling and parallel subgraph execution, each task completed in under 10 minutes—down from nearly 40 minutes on the baseline—enabling rapid, large-scale data generation.
@@ 5 Conclusion
We presented GraSP, a modular framework for synthetic data generation using graph-based, prompt-centric workflows. GraSP offers scalable, reproducible pipelines for language model training, featuring a low-code YAML interface, reusable subgraphs, agent nodes, and HuggingFace-native I/O. Its design supports diverse workflows, uniquely enabling multimodal inputs, subgraph reuse, conditional routing, and schema validation.
@@
GraSP accelerates dataset creation and promotes transparency and reuse in LLM development. Ongoing efforts must address risks like “model collapse” through mixed datasets and continuous quality control, ensuring GraSP’s utility across generative AI applications.
@@ A.1.2 Transformations
The RenameFieldsTransform is a lightweight transformation utility used in the GraSP pipeline to rename one or more fields in each record of the dataset. This is particularly useful for ensuring consistency in variable naming, aligning raw data to prompt-ready formats, or preparing input fields for downstream processing.
@@ A.1.3 Data Less Mode
In data-less mode, GraSP operates without any input source. Instead, it directly executes the graph and writes outputs based solely on intermediate or generated values. This is especially useful for bootstrapping datasets, performing zero-shot synthesis, or generating instructional data.
@@
GraSP supports both declarative and programmatic schema enforcement:
@@ A.5 Post-Generation Extensions
Appendix B Example GraSP YAML Configurations
This appendix provides example YAML configurations illustrating how GraSP pipelines are defined and composed using the data_config, graph_config, output_config, and schema_config sections. These examples demonstrate GraSP’s flexibility for data-driven and zero-shot pipelines, LLM orchestration, and safe output generation.

### v2

# Quality Tagging, and Management of Synthetic Data

@@ Abstract
The advancement of large language models (LLMs) is critically dependent on the availability of high-quality datasets for Supervised Fine-Tuning (SFT), alignment tasks like Direct Preference Optimization (DPO), etc. In this work, we present a comprehensive synthetic data generation framework, SyGra – that facilitates scalable, configurable, and high-fidelity generation of synthetic data tailored for these training paradigms. Our approach employs a modular and configuration-based pipeline capable of modeling complex dialogue flows with minimal manual intervention. This framework uses a dual-stage quality tagging mechanism, combining heuristic rules and LLM-based evaluations, to automatically filter and score data extracted from OASST-formatted conversations, ensuring the curation of high-quality dialogue samples. The resulting datasets are structured under a flexible schema, enabling seamless integration into diverse training workflows. Together, these innovations offer a robust solution for generating and managing synthetic conversational data at scale, significantly reducing the overhead of data preparation in LLM training pipelines. Our code and documentation are available at https://github.com/ServiceNow/SyGra
@@ 1 Introduction
To address this, we introduce SyGra (Graph-oriented Synthetic-data Pipeline), a general-purpose framework for scalable synthetic data generation. SyGra combines low-code, YAML-based configuration with modular, graph-driven orchestration to support complex workflows with branching, looping, and conditionals. It enables the reuse of graphs as subgraphs, ensures reliable execution through integrated validation and checkpointing, and natively supports multimodal inputs and agent based data generation. Additionally, SyGra offers unified dataset I/O across HuggingFace and local formats, supports quality tagging, and produces outputs compatible with OASST-style formatting for seamless downstream use.
Figure 1: High-level SyGra architecture.
@@ 2 Related Work
In summary, while existing frameworks each offer valuable features for synthetic data generation, they typically address isolated aspects of the broader workflow. SyGra stands out by providing a unified, extensible approach that brings together the critical capabilities needed for modern, complex, and multimodal data generation pipelines.
3 SyGra Framework
Table 1: Comparison of SyGra with popular frameworks across key capabilities.
Category Feature SyGra Distilabel SDG Curator Synthetic Data Kit
@@ 3 SyGra Framework
SyGra is a modular and extensible system designed for large-scale, programmable data generation. It supports configurable orchestration through a graph abstraction that enables reusable, auditable, and resumable workflows. The framework is designed for both research and production pipelines, with pluggable model backends and modular task authoring support.
@@ 3.1 System Architecture
SyGra is guided by three principles—Scalability (streaming data sources, resumable jobs, JSONL/Parquet/HF outputs), Modularity (YAML-defined DAG workflows with conditional logic), and Reusability (versioned, reusable graphs, nodes, and validators). Figure 1 shows its core components:
@@ 3.2 Pipeline Components
SyGra pipelines are defined declaratively in YAML, promoting low-code, reproducible workflow construction. Each pipeline consists of three configuration blocks:
@@
Specifies input and output sources, format handling (CSV, JSONL, Parquet), streaming options, and inline preprocessing (e.g., renaming, filtering, and combining). Supports both data-backed and data-less generation scenarios.
@@
Defines a DAG of computational nodes (Figure 2), each node can be configured to call LLMs, Python functions, agents, or subgraphs. Various node types are supported, such as:
  • llm: When a model needs to be called, we can use a LLM node model properties and role-based prompts, along with pre/post processors in Python code.
  • multi_llm: When we need to generate data at scale, we can use a multi-LLM node which allows configuration of load balanced model inferences between multiple endpoints.
  • lambda: To process the data during execution, we can utilize lambda nodes, which are mapped to Python functions.
  • agent: To perform end-to-end agentic behaviour, we can use agent nodes along with tools, which can be custom or Langchain tools.
  • subgraph: Complex flows can be splitted into smaller graphs i.e. subgraphs which can be reused inside the graph.
Once the nodes are defined, we connect them via egdes (Figure 3). Two types are supported – simple edge and conditional edge. Conditional edges are useful to build if-else flow and loops in the graph based on a condition written as a Python code.
@@
Finally, the graph is validated and compiled into a LangGraph-compatible representation. Refer to Appendix A for detailed configuration options and schema definitions.
Figure 2: Graph nodes
Figure 3: Graph edges
@@ 3.3 Key Features
SyGra brings together robust design abstractions and practical scalability for real-world use cases. Specifically, our contributions include:
@@
    Low-Code, Modular Graph Configuration: SyGra combines a YAML-based interface with LangGraph-style agents and a custom DAG engine, enabling concise, extensible definitions of complex workflows with branching, looping, and conditionals. B
@@
    Reusable Recipes (Subgraphs): This feature enables us to use common graph components which can be reused across tasks, promoting modularity. For instance, the Evolve INSTRUCT recipe (Figure  4) encapsulates a modular subgraph that receives seed instructions and applies either depth-based or breadth-based evolution strategies via a routing node (Strategy) [7]. This subgraph can be invoked repeatedly across different flows, enhancing composability and reducing redundancy.
    Figure 4: Instruction evolution subgraph and judgment loop used within SyGra pipelines.
@@
    Multimodal Support: SyGra extends beyond text-only workflows by natively handling audio and image inputs alongside text. Through unified I/O adapters, it transparently loads local or remote media in various formats, encodes them as base64 data URLs for LLM API compatibility, and supports multiple media fields per record. This enables workflows for tasks such as speech recognition, audio classification, document analysis, and visual QA. Round-tripping ensures outputs can be saved back into HuggingFace datasets in their original formats for reproducibility and downstream use.
@@
    Agentic Execution: SyGra enables the creation of autonomous, tool-using agents built on the ReAct reasoning-and-acting paradigm via LangGraph. Agent nodes extend LLM nodes with capabilities for dynamic tool invocation, multi-turn reasoning, and conditional decision-making. Developers can specify a library of callable tools, inject context-specific system messages at arbitrary conversation turns, and configure pre/post-processing hooks for fine-grained control over input and output. This allows pipelines to handle exploratory tasks, iterative search, and interactive decision flows in a modular, low-code manner.
@@
    Structured Output Generation: SyGra provides a flexible framework for generating and validating structured outputs from LLMs, reducing post-processing effort and ensuring reliable formats. It supports both class-based schemas (via Pydantic) and YAML-defined schemas, with automatic type handling and optional custom validation rules. Structured output generation works natively with OpenAI and vLLM models, and falls back to JSON schema validation for other backends. This allows developers to define precise field types, attach descriptions, and enforce constraints directly at generation time.
@@
    Resumability: SyGra supports fault-tolerant, restartable execution of long-running jobs. In the event of a failure, execution can gracefully shut down and later resume from the last recorded checkpoint without reprocessing completed steps. This is particularly valuable for large-scale or streaming workloads where partial progress should be preserved. Checkpoints store both intermediate outputs and node-level metadata, enabling accurate restoration of execution state.
@@ 4 Results and Impact
4.1 Experimental Setup
The evaluation was run on an 8-core CPU machine with 16 GB RAM using the SyGra framework. Model endpoints were deployed separately on vLLM with the Qwen 3 32B Instruct model, configured to use two GPUs with tensor parallelism, moderate CPU and memory resources, and optimizations such as chunked prefill with capped GPU memory usage.
The workload consisted of 10,000 input records, repeated across three trials at each concurrency level. To keep the focus on system behavior, the workflow was deliberately simple: a weighted sampler node selected tone and persona values, and an LLM node rephrased the input text using fixed generation parameters (deterministic outputs, max length 500 tokens). This design provided a controlled inference workload for measuring concurrency scaling without pipeline complexity.
4.2 Results
Performance was measured in terms of total wall-clock time required to process the workload under varying concurrency levels (denoted here as batch size).
Figure 6: Total completion time (seconds) for varying concurrency levels with one endpoint.
Three key patterns emerge:
  • Reduced completion time with higher concurrency. As concurrency increases from 10 to $\sim$500, total completion time decreases sharply, illustrating the efficiency of SyGra’s asynchronous execution.
  • Sustained efficiency at scale. Beyond $\sim$500 concurrent requests, completion time stabilizes around 900–1000 seconds, showing resilience under heavy load.
  • Server saturation at very high concurrency. At batch size 5000, total time increases slightly. This reflects a limitation of the underlying model server rather than SyGra itself. Users can mitigate this by adding multiple endpoints or increasing the computational resources allocated to the hosted model.
Note on Task Complexity. The experiment used a deliberately simple two-node workflow (sampling + rephrasing) to isolate concurrency effects. In this setting, a single vLLM endpoint was sufficient to handle high concurrency. However, in more complex tasks with multiple nodes and richer pipelines, SyGra exhibits stronger scaling when additional endpoints or resources are available, as state-of-the-art inference servers such as vLLM are capable of sustaining concurrent requests efficiently across distributed deployments.
4.3 Impact
These results underscore SyGra’s ability to deliver stable, predictable performance under extreme concurrency. Whereas traditional multi-threaded or multi-process inference servers often degrade sharply once concurrency exceeds a few hundred requests, SyGra sustains bounded completion times even at >>5000 concurrent requests.
The implications are significant:
  • Scalable deployments. SyGra enables operators to support thousands of simultaneous queries while maintaining consistent workload completion times.
  • Configurable performance. If saturation is observed at very high concurrency, performance can be extended by scaling to multiple endpoints or allocating additional resources to the hosted model.
  • Future readiness. As LLM-based systems (e.g., conversational agents, retrieval-augmented generation) face growing concurrency demands, SyGra provides a reliable inference layer that avoids bottlenecks at production scale.
In summary, SyGra demonstrates robust concurrency scaling, endpoint-agnostic performance, and stable execution times, making it a strong candidate for next-generation large-scale data syntheses.
5 Availability
SyGra is released as an open-source Python package and framework on PyPI555pip package: https://pypi.org/project/sygra/ (pip install sygra) and GitHub [^6] https://github.com/ServiceNow/SyGra.
@@ 6 Conclusion
We presented SyGra, a modular framework for synthetic data generation using graph-based, prompt-centric workflows. SyGra offers scalable, reproducible pipelines for language model training, featuring a low-code YAML interface, reusable subgraphs, agent nodes, and HuggingFace-native I/O. Its design supports diverse workflows, uniquely enabling multimodal inputs, subgraph reuse, conditional routing, and schema validation.
@@
SyGra accelerates dataset creation and promotes transparency and reuse in LLM development. Ongoing efforts must address risks like “model collapse” through mixed datasets and continuous quality control, ensuring SyGra’s utility across generative AI applications.
@@ A.1.2 Transformations
The RenameFieldsTransform is a lightweight transformation utility used in the SyGra pipeline to rename one or more fields in each record of the dataset. This is particularly useful for ensuring consistency in variable naming, aligning raw data to prompt-ready formats, or preparing input fields for downstream processing.
@@ A.1.3 Data Less Mode
In data-less mode, SyGra operates without any input source. Instead, it directly executes the graph and writes outputs based solely on intermediate or generated values. This is especially useful for bootstrapping datasets, performing zero-shot synthesis, or generating instructional data.
@@
SyGra supports both declarative and programmatic schema enforcement:
@@ A.5 Post-Generation Extensions
Appendix B Example SyGra YAML Configurations
This appendix provides example YAML configurations illustrating how SyGra pipelines are defined and composed using the data_config, graph_config, output_config, and schema_config sections. These examples demonstrate SyGra’s flexibility for data-driven and zero-shot pipelines, LLM orchestration, and safe output generation.